# Adanonymizer: Interactively Navigating and Balancing the Duality of Privacy and Output Performance in Human-LLM Interaction

### v1

@@ 4.1. Design Goals
Based on the previous work (Lison et al., 2021) and aiming to achieve the balance between privacy protection and model’s performance, the design goals were proposed as follows:
  1. (1)
@@ 4.3. Interaction Flow and Interface
The interface of Adanonymizer features a simple and efficient design (Kim and Lee, 2009), functioning as a floating plugin for the input text, making it adaptable to text-based chat and search platforms. Upon selecting the input text, users could activate Adanonymizer, with the selected text as the original input. Adanonymizer consists of three parts (see Figure 1): a control panel for proactive control through dragging (see Figure 1 (1)), an edit box for showing and editing the output (with changes highlighted, see Figure 1 (2)), and a click button to replace the original input (see Figure 1 (3)). The control panel provides an intuitive way for users to control privacy settings and model output performance levels, similar to a color selection tool. The privacy protection and model output performance levels are marked along the two axes of the panel. Users can adjust the privacy protection and model performances levels through clicking on the 2D palette. Based on these settings, Adanonymizer generates an anonymized version of the text, with variations depending on the level of anonymization users selected, and adopting to different phrases. The edit box features minimal formattign to reduce cognitive load and facilitate seamless system integration, a common design choice in previous literature (e.g., human-AI collaborative writing (Lee et al., 2024b, 2022)). Users could click on the “See Label” button to view the categories of the changed text, which were also highlighted. Users could also directly edit in the output editing box. When users click the “replace text” button, the original text is replaced by the anonymized version.
@@
1. Dragging the first progress bar restricts the range of the second.
2. A point on the single bar indicates different trade-offs between privacy and performance.
3. In the two-dimensional palette, the upper-right points are unreachable, showing that privacy protection and model performance cannot be achieved simultaneously.
@@ 5. Implementation of Adanonymizer
Adanonymizer is implemented in the form of a browser plug-in to integrate with different LLMs. The implementation used Python and Javascript. The backend algorithms utilized the GPT-4o API (API version GPT-4o-2024-05-13). In this section, we first outline its framework, followed by a detailed description of the anonymization process.
@@ 5.1. System Structure
The system contained the panel for users to control, the mapping from users’ control to the selective anonymization strategy (denoted as mapping strategy) and the back-end anonymization algorithm (denoted as anonymization algorithm). The latency of the whole system was less than ten seconds, as described in Section 6.4, which did not affect participants’ experience.
@@ 6.1. Participants and Apparatus
We recruited 36 participants (21 males, 15 females, aged 18 to 26, M=21.8, SD=1.6) in China from the XX campus (anonymized for submission) through snowball sampling (Goodman, 1961). All participants were familiar with interacting with LLMs and had a self-reported LLM usage experience of 8.2 months (SD=2.4). However, they were not familiar with the technical perspective of privacy preservation. No participant dropped the experiment and all participants received 150 RMB as compensation. The experiment was conducted through an online environment provided as a web service developed by the experimenter through Flask. We implemented the experiment platform using the API of GPT-4o, while we implemented the anonymizers in the form of pop-ups based on the experiment platform.
@@ 6.2. Experiment Design
We conducted a one-factor within-subjects study with the anonymization technique as the only factor. Adanonymizer and its ablation techniques operate at three different interactive levels. We set three levels as three different techniques during comparison: No interactive (denoted as Automatic), Single interactive (denoted as Privacy Only) and Full interactive (Adanonymizer). In no interactive, the system would modify all the information automatically without user input. In “Privacy Only” setting, the user could control the privacy protection level and the system would control which information to modify under the specific privacy protection level. It always maximize utility during modification. In Adanonymizer setting, the user should control both the privacy protection level and model performance preservation level. Notably, the No Interaction and Single Interaction settings serve as ablation comparisons to the Full Interaction setting.
@@ 7.3. Design Implications
Incorporating user perspective directly into privacy preserving technique is possible. Incorporating user perspectives into privacy-preserving techniques is feasible, as demonstrated by the successful implementation of Adanonymizer, which uses subjective ratings. Based on Privacy Calculus theory (Laufer and Wolfe, 1977), user ratings can balance privacy and utility in protecting image and voice data by selectively masking or revealing items and regions based on user concerns and preferences (Hasan et al., 2019; Wang et al., 2023a). For example, items of higher importance should remain unmasked, while sensitive but less important ones should be concealed. This approach can also be applied to smartphones (Shih et al., 2015) and smart homes (Chi et al., [n. d.]), where data with lower privacy concerns and higher preferences is transmitted, while highly sensitive data is blocked.
Interaction designs could leverage interactive 2D panels for balancing trade-offs such as privacy protection and model performance. Adanonymizer enables users to actively manage the privacy-utility trade-off through an intuitive interface, allowing for informed decisions about data usage. Its modular design adapts to varying privacy risk levels, offering scalability across applications like image and video privacy protection (Wang et al., 2023a). Users can selectively adjust the interface to achieve their desired balance of privacy protection and image recognition performance.
Privacy protection methods can enhance their transparency through metaphors like controllable panel. In this paper, we utilize intuitive panels to convey and manage privacy protection settings for pseudo-anonymization. User control determined the privacy protection level, which governed the background anonymization algorithm. Many algorithms, such as DP-based methods (Xiong et al., 2020; Franzen et al., 2022), are inherently opaque and require clear communication of their parameters, settings, and effects to users. Our approach to visualizing privacy protection and model performance can be extended to other methods, such as DP (Fernandes et al., 2019) or encryption (Lin et al., 2024a), by illustrating $\epsilon$ in DP and cryptographic strength in encryption algorithms.
@@ 9. Limitation and Future Works
Third, we only selected the privacy risk towards re-training (Zhang et al., 2024a) to study. The most significant difference of human-LLM interaction the traditional online media interaction (Bryce and Fraser, 2014) and human-AI interaction (Amershi et al., 2019) was that LLMs frequently needed novel data for fine-tuning and model improvement. Although there exist other data misuse and privacy risk beyond re-training, different usage exhibit different privacy risk patterns for participants, which were out of this paper’s scope. We deemed those as our future work.
Moreover, we utilized large models for Adanonymizer to avoid using users’ private data. In future deployments, companies could train models with compliant data to achieve the same functionality.

### v2

@@ 4.1. Design Goals
Building on prior work (Lison et al., 2021) and aiming to balance privacy protection and model performance, we introduced the following design goals:
@@ 4.2. Proactiveness of Adanonymizer
4.3. Palette Design
For the 2D palette, it has different layout manners and mapping strategies to balance privacy and utility. Regarding the dimension of layouts, the palette could adopt a discrete layout or continuous layout. Discrete layout has the advantages of granting users easier control through choosing from a few options, while continuous layout could guarantee users a more smooth journey through tuning the optimal points, while in the same time observing the change of output in real time, as shown in Figure 5.
(a) Discrete.
(b) Continuous with magnet effect.
(c) Continuous.
Figure 5. Different layout designs of the 2D palette.
We evaluated several strategies for mapping privacy and utility. One option involved directly mapping the values onto the 2D palette, either linearly or with equal intervals. Linear mapping assigns equal intervals to privacy items with similar sensitivity scores, reflecting users’ perceived differences. In contrast, equal interval mapping assigns the same spacing regardless of sensitivity, simplifying item selection. Alternatively, piecewise mapping could be used, where high-privacy and high-utility items are assigned larger intervals than low-privacy and low-utility ones, reflecting participants’ priorities. Based on a pilot study, we chose linear mapping, as participants preferred its transparency, which enhanced their sense of control.
@@ 4.4. Interaction Flow and Interface
The interface of Adanonymizer features a simple and efficient design (Kim and Lee, 2009), functioning as a floating plugin for the input text, making it adaptable to text-based chat and search platforms. Upon selecting the input text, users could activate Adanonymizer, with the selected text as the original input. Adanonymizer consists of three parts (see Figure 1): a control panel for proactive control through dragging (see Figure 1 (1)), an edit box for showing and editing the output (with changes highlighted, see Figure 1 (2)), and a click button to replace the original input (see Figure 1 (3)). The control panel provides an intuitive way for users to control privacy settings and model output performance levels, similar to a color selection tool. The privacy protection and model output performance levels are marked along the two axes of the panel. Users can adjust the privacy protection and model performances levels through clicking on the 2D palette. Based on these settings, Adanonymizer generates an anonymized version of the text, with variations depending on the level of anonymization users selected, and adopting to different phrases. The edit box features minimal formatting to reduce cognitive load and facilitate seamless system integration, a common design choice in previous literature (e.g., human-AI collaborative writing (Lee et al., 2024b, 2022)). Users could click on the “See Label” button to view the categories of the changed text, which were also highlighted. Users could also directly edit in the output editing box. When users click the “replace text” button, the original text is replaced by the anonymized version.
@@
1. (Multi-constraint sliders) Dragging the first progress bar restricts the range of the second (Figure 6a).
2. (Single slider) A point on the single bar indicates different trade-offs between privacy and performance (Figure 6b).
3. (Panel) In the two-dimensional palette, the upper-right points are unreachable, showing that privacy protection and model performance cannot be achieved simultaneously (Figure 6c).
(a) Multi-constraint sliders.
(b) Single slider.
(c) Panel.
Figure 6. The candidate interface designs for privacy-utility balancing.
@@ 5. Implementation of Adanonymizer
Adanonymizer is implemented in the form of a local plugin to prevent privacy and ethical issues. We used qwen-2.5-7b-instruct for implementation although the backbone model could potentially be transfered to other architectures. The implementation used Python and Javascript. In this section, we first outline its framework, followed by a detailed description of the anonymization process.
@@ 5.1. System Structure
The system contained the panel for users to control, the mapping from users’ control to the selective anonymization strategy (denoted as mapping strategy) and the back-end anonymization algorithm (denoted as anonymization algorithm).
@@ 5.2. Mapping Strategy
We further normalized the score to $[0,1]$ through first subtracting the minimal value and then divide the range. After normalization, the whole information space would be mapped to the two-dimensional panel and participants could have full control.
@@ 5.3. Anonymization Algorithm
To instantiate the recognition process locally, we adopted a qwen-2.5-7b-instruct because of its superior performance on open-source benchmarks. To make the processing possible on personal laptops, we further adopted quantization and runtime optimization techniques, involving using int8 quantization.
@@ 6.1. Participants and Apparatus
We recruited 36 participants (21 males, 15 females, aged 18 to 26, M=21.8, SD=1.6) in China from the XX campus (anonymized for submission) through snowball sampling (Goodman, 1961). All participants were familiar with interacting with LLMs and had a self-reported LLM usage experience of 8.2 months (SD=2.4). However, they were not familiar with the technical perspective of privacy preservation. No participant dropped the experiment and each participant received 150 RMB as the compensation. The experiment was conducted through an online environment provided as a web service developed by the experimenter through Flask. We implemented the experiment platform using the API of GPT-4o, while we implemented the anonymizers in the form of pop-ups based on the experiment platform.
@@ 6.2. Experiment Design
We conducted a two-factor within-subjects study with the anonymization technique and task as two within factors. Adanonymizer and its ablation techniques operate at three different interactive levels. We set three levels as three different techniques during comparison: No interactive (denoted as Automatic), Single interactive (denoted as Privacy Only) and Full interactive (Adanonymizer). In no interactive, the system would modify all the information automatically without user input. In “Privacy Only” setting, the user could control the privacy protection level and the system would control which information to modify under the specific privacy protection level. It always maximize utility during modification. In Adanonymizer setting, the user should control both the privacy protection level and model performance preservation level. Notably, the No Interaction and Single Interaction settings serve as ablation comparisons to the Full Interaction setting.
@@ 6.4.4. Subjective Comments
6.4.5. Trade-off Between Privacy Anonymization and Utility
We interviewed participants to understand their trade-off behavior between privacy anonymization and performance.
Control Path
We summarized three themes and strategies for users’ control.
Tuning along the line. 15 participants followed the highlighted balance line for tuning their control. “I would first take notice of the line and then select points near the line to reach the optimal balance point for me.” (P10).
Target for balance than extreme. 13 participants first searched and stayed on the upper right part of the panel, which indicates the balance between privacy and utility. They thought “These points were more suitable because they guaranteed privacy and utility to a great extent in the same time instead of only focusing on one.”
Focus on angles which indicate Optimized Points. 20/36 participants always seek for the corners of the polyline rather than other points. Theoretically the optimal convex line should be consecutive along the privacy-utility panel. However, because the discrete nature of text, anonymization and privacy categories, the curve is a polyline rather than a smooth curve.
Final Balance State
We found most of the time all participants would achieve the balance state of privacy and utility. However, nuanced patterns exist and themes were summarized below:
Participants select different balance state for different tasks. For study-related tasks 20/36 participants disclose more information and thus select privacy protection higher, however 10/36 participants cared more about the task performance in work and life-related tasks, thus resulting in more task-oriented privacy protection settings.
Participants select different balance state based on different personalities. A few participants (5/36) cared extremely about their privacy and exhibit extreme patterns for some specific tasks, resulting selecting to anonymize all text. Interestingly, no participant only cared about task performance without considering privacy. All participants’ selection were near the line or on the line, indicating that participants always sought for the best balance between privacy and utility.
@@ 7.3. Design Implications
Enabling users to proactively control the privacy-utility balancing based on privacy risk and utility modeling. Based on privacy calculus theory (Laufer and Wolfe, 1977), application fields such as healthcare (Ragan et al., 2018) could leverage the modeling framework to engage users in proactively control their trade-off between privacy and utility, thus enhancing the automatic algorithmic-based balancing methods.
Interaction designs could leverage interactive 2D panels for balancing trade-offs such as privacy protection and model performance.
Future interaction design for multimodal frameworks (Hu et al., 2024) or privacy protection for visual language models (Wang et al., 2023a) could leverage similar framework to manage the privacy-utility trade-off through the 2D panels, enhancing transparency and ease-of-control.
Privacy protection methods can enhance their transparency through controllable panel metaphors. Many algorithms, such as DP-based methods (Xiong et al., 2020; Franzen et al., 2022), are inherently opaque and require clear communication of their parameters, settings, and effects to users. The panel based approach to visualize privacy protection and model performance can also be extended to other methods, such as DP (Fernandes et al., 2019) or encryption (Lin et al., 2024a), by illustrating $\epsilon$ in DP and cryptographic strength in encryption algorithms, where user control could determine the privacy budget and govern the background protection algorithm.
@@ 9. Limitation and Future Works
Third, we only selected the privacy risk towards re-training (Zhang et al., 2024a) to study. The most significant difference of human-LLM interaction the traditional online media interaction (Bryce and Fraser, 2014) and human-AI interaction (Amershi et al., 2019) was that LLMs frequently needed novel data for fine-tuning and model improvement. Although there exist other data misuse and privacy risk beyond re-training, different usage exhibit different privacy risk patterns for participants, which were out of this paper’s scope. We deemed those as our future work. Moreover, we utilized large models for Adanonymizer to avoid using users’ private data. In future deployments, companies could train models with compliant data to achieve the same functionality.